a 5L sample of CO2 at 540torr and 25 degrees celsius expands to 15L at 32 degrees Celsius. What is the pressure of the new container?
1
Expert's answer
2018-03-07T12:01:08-0500
To solve this type of problem, we use the combined gas law
(p_0×V_0)/T_0 = (p_1×V_1)/T_1
It is better to translate the temperature into Kelvin
T=t^° C+273
T_1= 25^(° )C+ 273=298 K
T_2=32^° C+273=305 K
p_1=(p_(0 )×V_(0 )×〖 T〗_1)/(T_0×V_1 ) =(540×5×305)/(298×15) =184.23 torr